Could someone help me with this problem? f(x)=(x-3)/(x+2). Thank you so very much!
Answer by Nate(3500)

in graghing:
f(x)=(x-3)/(x+2)
is the same as:
y = (x-3)/(x+2)
it would be shaped as a sideways hyperbola....
----------------------------------------------------------------------
in solving:
pick a point such as x=3
f(x)=(x-3)/(x+2)
f(3)=(3-3)/(5) fill in 3 for the x's
f(3)=0
REMEMBER that x cannot be -2 for that is undefined!